OpenStudy (anonymous):

Paul is painting a square wall with an area of 115 square feet. To the nearest foot, what is the length of the wall?

OpenStudy (p0sitr0n):

A = c*c 115ft^2 = C^2 \[\sqrt{115} = c\] c = 11 ft